# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

### v1

@@ Abstract
Large recommender models have extended LLMs as powerful recommenders via encoding or item generation, and recent breakthroughs in LLM reasoning synchronously motivate the exploration of reasoning in recommendation. Current studies usually position LLMs as external reasoning modules to yield auxiliary thought for augmenting conventional recommendation pipelines. However, such decoupled designs are limited in significant resource cost and suboptimal joint optimization. To address these issues, we propose $\text{R}^{2}\text{ec}$, a unified large recommender model with intrinsic reasoning capabilities. Initially, we reconceptualize the model architecture to facilitate interleaved reasoning and recommendation in the autoregressive process. Subsequently, we propose RecPO, a corresponding reinforcement learning framework that optimizes $\text{R}^{2}\text{ec}$ both the reasoning and recommendation capabilities simultaneously in a single policy update; RecPO introduces a fused reward scheme that solely leverages recommendation labels to simulate the reasoning capability, eliminating dependency on specialized reasoning annotations. Experiments on three datasets with various baselines verify the effectiveness of $\text{R}^{2}\text{ec}$, showing relative improvements of 68.67% in Hit@5 and 45.21% in NDCG@20. Code available at https://github.com/YRYangang/RRec.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in contextual understanding and open-ended generation [1, 2, 3, 4]. This success has catalyzed the development of large recommender models that inherit and specialize in the advantages of LLMs for recommendation. Current approaches can be divided into two main streams: one employs LLMs as powerful encoders to embed users (their historical interactions) [5, 6, 7], while the other reformulates item prediction into the autoregressive generation of item identifiers [4, 8]. These large recommender models exhibit remarkable generalization capabilities, achieving advanced performance across diverse application scenarios, including cold-start recommendations [9], cross-domain personalization [10], and long-tail item prediction [11, 12].
@@
Existing studies have preliminarily explored LLM reasoning for recommendation, including user preference analysis [15, 16], synthetic item profile enrichment [17, 18], user search query rewriting [19, 20], and rating prediction [21]. These approaches typically take LLMs as additional reasoning modules to augment the original recommendation pipeline, decoupling reasoning from item prediction. However, such designs introduce crucial limitations: 1) Significant resource cost. Two distinct modules, namely the large reasoning model and the recommendation model, must be trained, checkpointed, and served in parallel, inflating both memory footprint and inference latency. 2) Suboptimal joint optimization. Reasoning and recommendation modules can only be trained by freezing one while updating the other—first update the reasoning module, then switch to the recommendation module, and continue alternating in this manner [15]. This alternate optimization scheme prevents gradient flow across the pipeline, precluding true end-to-end learning, hindering fine-grained alignment between reasoning and ranking objectives, and ultimately leading to suboptimal convergence. In this work, we aim to develop a unified large recommender model with intrinsic reasoning capabilities, exploring new technical solutions for reasoning-enhanced recommendation systems.
It is non-trivial to develop a unified large recommender model with reasoning due to the following aspects: 1) Model design. Recommendation is traditionally framed as a discriminative ranking task, whereas reasoning is a generative process. How to incorporate both generative reasoning and discriminative recommendation into one single model architecture is worth careful consideration. 2) Training optimization. It is also challenging to stimulate and optimize the reasoning capabilities of large recommender models. On the one hand, recommendation domains lack annotated reasoning traces, as personalized recommendation rationale is hardly obtainable, making supervised learning impractical. On the other hand, Reinforcement Learning (RL) [20, 22, 23] is a natural alternative but encounters unique problems in reward and training objective design within the recommendation task. Since reasoning and recommendation capabilities interact within a single model, this necessitates a training framework that effectively bridges these two aspects.
To address the above issues, we propose a unified large Recommender model with intrinsic Reasoning, dubbed as $\text{R}^{2}\text{ec}$. 1) In the model design, as illustrated in Figure 1, we introduce the recommendation head to the LLM backbone, alongside the original language-modeling head. This setup allows the model to first generate reasoning tokens autoregressively and then predict the item in the last step. This model architecture integrates both generation and recommendation tasks, while the next crucial step is to simulate the reasoning capabilities tailored for accurate item predictions. 2) In training optimization, we propose an RL-based training framework without human reasoning annotations. First, we sample trajectories from the model, which essentially consist of reasoning-then-item sequences, similar to the reasoning-then-answer sequences in LLMs. Second, unlike Deepseek-R1, which defines rewards based on answer accuracy, we found it insufficient to set the reward solely using item recommendation metrics. Therefore, we propose a fused reward scheme that combines discrete ranking rewards with continuous similarity rewards. With the sampled trajectories and assigned rewards, a joint reasoning-and-recommendation training objective is introduced to optimize the model for improving the recommendation through reasoning.
We conducted extensive experiments to compare $\text{R}^{2}\text{ec}$ with various traditional, LLM-based, and reasoning-augmented recommendation baselines. The experimental results on three datasets illustrate $\text{R}^{2}\text{ec}$ significantly outperforms all baselines with relative improvements of 68.67% in Hit@5 and 45.21% in NDCG@20, verifying the effectiveness of the unified reasoning-recommendation model design. We also performed the ablation study and 5 in-depth analysis experiments, demonstrating that $\text{R}^{2}\text{ec}$ possesses strong generalization capabilities with different optimization algorithms. More surprisingly, we observed a similar phenomenon in LLM reasoning that the reasoning length increases as training progresses.
The key contributions in this work are summarized as follows:
{adjustwidth}
[…]
2 Related Work
2.1 Reinforcement Learning for LLM Reasoning
OpenAI’s GPT-o1 demonstrated that scaling reinforcement learning (RL) training with large compute budgets enables models to internalize reasoning as a differentiable policy, achieving state-of-the-art performance in emergent meta-reasoning tasks [24]. Early RLHF methods trained a reward model on human preference data and fine-tuned the policy using Proximal Policy Optimization (PPO) [23]; however, PPO’s multiple optimization passes introduce implementation complexity. To streamline RL tuning, Direct Preference Optimization (DPO) [25] performs single-step policy updates, trading simplicity for potential off-policy bias. Alternative estimators, such as Group Relative Policy Optimization (GRPO) [22], derive baselines from group-level rewards, eliminating the need for a separate critic, while REINFORCE’s leave-one-out (RLOO) [26] computes advantages by excluding each trajectory sequentially. Building on these advancements, DeepSeek-Zero [27] removes reliance on supervised fine-tuning by driving reasoning emergence purely through intrinsic RL rewards, and DeepSeek-R1 further integrates rule-driven reward engineering with self-play verification to enhance reasoning robustness.
2.2 Large Language Model-based Recommendation
Recent advances in LLM-based recommendation typically follow two paradigms: 1) LLM-enhanced, where LLMs enrich recommendation pipelines with additional features, and 2) LLM-centric, where recommendation is framed as a generative task via item identifiers [1, 2, 3, 4, 28, 29], or LLMs serve as encoders to embed users and items [6, 30]. We term models under the second paradigm as large recommender models. In the first paradigm, LLM-generated embeddings are utilized to enhance conventional recommenders. Adoption of LLM-generated text are not new, typical examples includes user intent [15, 31], search queries [20], item summaries [18], or rationales [32]. Among them, SLIM [32], LangPTune [15], and Rec-R1 [20] further optimize LLM for finer rationale generation with RL. However, such designs introduce high resource demands for training and serving both models, and the connection between two modules remains gradient-less, blocking true end-to-end learning, resulting in suboptimal convergence and performance. Large recommender models typically recommend through autoregressive generation of item identifiers [4, 28] or titles [1, 2, 3, 29], or by embedding user interaction sequences [6, 30]. Large recommender model has demonstrated remarkable potential, but current methods fail to utilize reasoning ability of LLMs. To the best of our knowledge, no existing large recommender model improves recommendation over explicit reasoning generation, which is the gap that motivated our work.
@@ 3.2 Reinforcement Learning for Large Reasoning Models
Sampling-based advantage estimation is widely adopted in recent advances [22, 26, 34, 35]; below we describe its basic pipeline and two typical computation methods. Given an input ${x}$, a group of $G$ different trajectories $\{{o}_{i}\}_{i=1}^{G}$, are sampled from $\pi_{\theta_{\text{old}}}$. Existing studies widely obtain these trajectories via top-$K$ sampling with temperature, where $\theta_{\text{old}}$ refers to the frozen policy parameters before current update. Each trajectory $o_{i}$ receives a scalar reward $R_{i}$, which will then be used to compute the trajectory-level advantages via two widely adopted approaches, namely GRPO [22] and RLOO [26]:
@@ 4.1 Model Design
Inference begins with feeding the tokenized user prompt ${x}_{u}$ (template of prompts can be found in Appendix E) into the transformer-based backbone, producing an initial hidden state $\bm{h}_{0}$. The language-modeling head then maps $\bm{h}_{0}$ to the first reasoning token $o_{1}$. This process continues autoregressively, yielding a sequence of $T$ reasoning tokens ${o}_{1:T}$. The final hidden state of the generated sequence $\bm{h}_{T}$ is then fed into the recommendation head, where each candidate item $v\in\mathcal{V}$ is scored by an inner product $s(v)={\bm{h}}_{T}^{\top}\mathbf{H}_{\mathcal{V}}[v],$ $v\in\mathcal{V},$ which determines the final ranking. This mechanism yields a tight reasoning–recommendation coupling, since both the language-modeling head and recommendation head share the same hidden-state space as input, reasoning directly reshapes $\bm{h}_{T}$ and thus yielding more accurate recommendation scores $s(v)$. Such alignment ensures that reasoning optimization (Section 4.2) contributes directly to finer recommendation.
@@ 4.2 Optimization: RecPO
Our goal is to train the policy $\pi_{\theta}$ to jointly performs reasoning and recommendation, i.e., it must generate coherent reasoning sequences to rank the target item accurately. Accordingly, we structure our optimization workflow in three parts. First, we introduce the trajectory sampling strategy that draws multiple reasoning trajectories for each user (Section 4.2.1). Second, we describe reward and advantage estimation, where discrete ranking signals and softmax similarities are fused into a single scalar award for the above sampled trajectories (Section 4.2.2). Finally, we formulate the training objective, which blends reasoning-level and recommendation-level updates through a clipped-ratio loss (Section 4.2.3). For a complete description of our end-to-end training and inference pipeline, please refer to Appendix D.
@@ 4.2.3 Training Objective
$\quad r_{i,t}(\theta)$ =$=\begin{cases}\frac{\pi_{\theta}(o_{i,t}|x_{u},{o}_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|x_{u},{o}_{i,<t})},&\text{if }t\leqslant T\text{ (reasoning)}\\ \frac{\pi_{\theta}(v^{+}|x_{u},{o}_{i,\leqslant T})}{\pi_{\theta_{\text{old}}}(v^{+}|x_{u},{o}_{i,\leqslant T})},&\text{if }t=T+1\text{ (recommendation)}.\end{cases}$ (6)
@@
In Eqn.(8), all trajectories contribute to token-level policy updates via the standard clipped objective $\ell_{\epsilon}\!\bigl{(}r_{i,t}(\theta),A_{i}\bigr{)}$. This design ensures the policy continues to learn from diverse reasoning behaviors. For the last recommendation action, only the trajectory with the highest advantage, identified by $i^{\star}=\arg\max_{j}A_{j}$, contributes gradients to recommendation optimization. This design concentrates recommendation learning with the most promising reasoning path, while the presence of other $G-1$ reasoning paths contributes to ample exploration in reasoning actions, thus preserving exploration and ensuring effective recommendation learning.
@@ 5.1 Setups
Following previous work on large recommender models [3, 2, 1, 29], we conducted experiments using real-world datasets sourced from Amazon, including CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments). Dataset statistics and preprocessing steps can be referred to Appendix A. We utilized two commonly used metrics: Hit Rate (H@K) and Normalized Discounted Cumulative Gain (N@K), with cutoff $K$ set to $5$, $10$ and $20$. We adopted the full set recommendation setting, where metrics are computed over the entire item set, providing a better reflection of practical scenarios.
@@
To comprehensively evaluate $\text{R}^{2}\text{ec}$, we selected competitive baselines from various categories, including traditional sequential recommenders (GRU4Rec [38], Caser [39], SASRec [40]), LLM-based recommender (TIGER [8], BigRec [2], $D^{3}$ [3]), and reasoning augmented recommendation systems (LangPTune [15]). More baseline details can be found in Appendix B. We used Qwen and Gemma as the LLM backbones. More implementation details are described in Appendix C.

### v3

# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1,4 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

@@ Abstract
Large recommender models have extended LLMs as powerful recommenders via encoding or item generation, and recent breakthroughs in LLM reasoning synchronously motivate the exploration of reasoning in recommendation. In this work, we propose $\text{R}^{2}\text{ec}$, a unified large recommender model with intrinsic reasoning capability. $\text{R}^{2}\text{ec}$ introduces a dual-head architecture that supports both reasoning chain generation and efficient item prediction in a single model, significantly reducing inference latency. To overcome the lack of annotated reasoning data, we design RecPO, a reinforcement learning framework that optimizes reasoning and recommendation jointly with a novel fused reward mechanism. Extensive experiments on three datasets demonstrate that $\text{R}^{2}\text{ec}$ outperforms traditional, LLM-based, and reasoning-augmented recommender baselines, while further analyses validate its competitive efficiency among conventional LLM-based recommender baselines and strong adaptability to diverse recommendation scenarios. Code and checkpoints available at https://github.com/YRYangang/RRec.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in contextual understanding and open-ended generation [1, 2, 3, 4, 5]. This success has catalyzed the development of large recommender models that inherit and specialize in the advantages of LLMs for recommendation. Current approaches can be divided into two main streams: one employs LLMs as powerful encoders to embed users (their historical interactions) [6, 7, 8], while the other reformulates item prediction into the autoregressive generation of item identifiers [4, 9]. These large recommender models exhibit remarkable generalization capabilities, achieving advanced performance across diverse application scenarios, including cold-start recommendations [10], cross-domain personalization [11], and long-tail item prediction [12, 13].
@@
Existing studies have preliminarily explored LLM reasoning for recommendation, including user preference analysis [16, 17], synthetic item profile enrichment [18, 19], user search query rewriting [20, 21], and rating prediction [22]. These approaches typically take LLMs as additional reasoning modules to augment the original recommendation pipeline, decoupling reasoning from item prediction. However, such designs introduce crucial limitations: 1) Significant resource cost. Two distinct modules, namely the large reasoning model and the recommendation model, must be trained, checkpointed, and served in parallel, inflating both memory footprint and inference latency. 2) Suboptimal joint optimization. Reasoning and recommendation modules can only be trained by freezing one while updating the other [16, 23]. This alternate optimization scheme prevents gradient flow across the pipeline, precluding true end‑to‑end learning, hindering fine‑grained alignment between reasoning and ranking objectives, and ultimately leading to suboptimal convergence. In this work, we aim to develop a unified large recommender model with intrinsic reasoning capabilities, exploring new technical solutions for reasoning-enhanced recommendation systems.
It is non-trivial to develop a unified large recommender model with reasoning due to the following aspects: 1) Model design. Most large recommender models rely on autoregressive decoding of item identifiers (IDs), which can be computationally expensive [1, 2, 3, 4, 24, 25, 26]. Introducing reasoning would inevitably exacerbate the latency. Thus, a key challenge is to integrate reasoning while maintaining acceptable inference speed. 2) Training optimization. A major obstacle in training the reasoning recommender is the scarcity of high-quality, annotated reasonings in recommendation domains. Unlike general QA tasks where solution rationales can be curated, the underlying rationale for recommendation is inherently subjective and difficult to collect at scale. Reinforcement Learning (RL) [21, 27, 28] is a natural alternative, but it introduces recommendation-specific challenges in reward specification and objective design. The training framework must effectively couple the reasoning and recommendation objectives to ensure co-optimization.
To address the above issues, we propose a unified large Recommender model with intrinsic Reasoning, dubbed as $\text{R}^{2}\text{ec}$. Our approach is built on two pillars: 1) Model design. As illustrated in Figure 1, we equip an LLM with a dual-head design: a standard language head for reasoning and a novel recommendation head for item prediction. The recommendation head maps the model’s output to a shared semantic item embedding space. This allows the model to first generate a reasoning chain autoregressively and then select the final item in a single step, replacing the slow auto-regressive decoding of item IDs with a more efficient "next-item prediction". 2) Optimization. we propose RecPO, an RL-based training framework without human reasoning annotations. First, for a given input, we sample multiple diverse reasoning paths, each culminating in an item recommendation, forming "reasoning-then-item" sequences analogous to those in reasoning LLMs. Second, we found it insufficient to set the reward naively as ranking metrics, we thus propose a fused reward scheme that combines discrete ranking rewards with continuous similarity rewards. Finally, we formulate a joint training objective that leverages the rewards to optimize the model, encouraging it to develop reasoning pathways that lead to high-quality recommendations.
We conducted extensive experiments to evaluate the efficacy of $\text{R}^{2}\text{ec}$ by comparing it with various traditional, LLM-based, and reasoning-augmented recommendation baselines. The experimental results on three datasets demonstrate that $\text{R}^{2}\text{ec}$ significantly outperforms all baseline methods, verifying the effectiveness of its unified architect and its ability to leverage reasoning for superior recommendation quality. Our comprehensive evaluation comprises nine analyses, including a detailed profiling of the model’s reasoning behavior which reveals its context-aware capabilities; an efficiency analysis showing competitive inference speed among LLM-based recommenders; ablation and optimization studies on key components; and additional investigations into scalability, generalization and case studies.
Key contributions are as follows:
  • We introduce $\text{R}^{2}\text{ec}$, the first unified large recommender model that intrinsically integrates reasoning and recommendation within a single architecture, resulting in significantly lower inference latency compared to both reasoning-augmented and conventional LLM-based recommender baselines.
  • We propose RecPO, an RL framework for recommendation that enables model optimization without reasoning annotations. We derive a joint optimization objective and propose a fused reward scheme, yielding performance that significantly surpasses all baselines.
  • Beyond metrics, we provide an in-depth analysis of the model’s reasoning behavior, demonstrating that $\text{R}^{2}\text{ec}$ automatically develops context-aware reasoning strategies tailored to specific domains and user contexts, offering meaningful interpretability for its recommendations.
@@ 2.2 Reinforcement Learning for Large Reasoning Models
Sampling-based advantage estimation is widely adopted in recent advances [27, 31, 32, 33]; below we describe its basic pipeline and two typical computation methods. Given an input x{x}, a group of $G$ different trajectories $\{{o}_{i}\}_{i=1}^{G}$, are sampled from $\pi_{\theta_{\text{old}}}$. Existing studies widely obtain these trajectories via top-$K$ sampling with temperature, where $\theta_{\text{old}}$ refers to the frozen policy parameters before current update. Each trajectory $o_{i}$ receives a scalar reward $R_{i}$, which will then be used to compute the trajectory-level advantages via two widely adopted approaches, namely GRPO [27] and RLOO [31]:
@@ 3.1 Model Design
Inference begins with feeding the tokenized user prompt ${x}_{u}$ (template of prompts can be found in Appendix G) into the transformer-based backbone, producing an initial hidden state $\bm{h}_{0}$. The language‐modeling head then maps $\bm{h}_{0}$ to the first reasoning token $o_{1}$. This process continues autoregressively, yielding a sequence of $T$ reasoning tokens ${o}_{1:T}$. The final hidden state of the generated sequence $\bm{h}_{T}$ is then fed into the recommendation head, where each candidate item $v\in\mathcal{V}$ is scored by an inner product $s(v)={\bm{h}}_{T}^{\top}\mathbf{H}_{\mathcal{V}}[v],$ $v\in\mathcal{V},$ which determines the final ranking. This mechanism yields a tight reasoning–recommendation coupling, since both the language-modeling head and recommendation head share the same hidden‐state space as input, reasoning directly reshapes $\bm{h}_{T}$ and thus yielding more accurate recommendation scores $s(v)$. Such alignment ensures that reasoning optimization (Section 3.2) contributes directly to finer recommendation.
@@ 3.2 Optimization: RecPO
The goal is to train the policy $\pi_{\theta}$ to jointly perform reasoning and recommendation, i.e., it must generate coherent reasoning sequences to rank the target item accurately. Accordingly, we structure our optimization workflow in three parts. 1) we introduce the trajectory sampling strategy that draws multiple reasoning trajectories for each user (Section 3.2.1). 2) we describe reward and advantage estimation, where discrete ranking signals and softmax similarities are fused into a single scalar award for the above sampled trajectories (Section 3.2.2). 3) we formulate the training objective, which blends reasoning-level and recommendation-level updates through a clipped-ratio loss (Section 3.2.3). Complete description of training and inference pipeline are in Appendix E.
@@ 3.2.3 Training Objective
$\quad r_{i,t}(\theta)$ $=\begin{cases}\frac{\pi_{\theta}(o_{i,t}|x_{u},{o}_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|x_{u},{o}_{i,<t})},&\text{if }t\leqslant T\text{ (reasoning)}\\ \frac{\pi_{\theta}(v^{+}|x_{u},{o}_{i,\leqslant T})}{\pi_{\theta_{\text{old}}}(v^{+}|x_{u},{o}_{i,\leqslant T})},&\text{if }t=T+1\text{ (recommendation)}.\end{cases}$ (6)
@@
In Eqn.(8), all trajectories contribute to token-level policy updates via the standard clipped objective $\ell_{\epsilon}\!\bigl(r_{i,t}(\theta),A_{i}\bigr)$. This design ensures the policy continues to learn from diverse reasoning behaviors. For the last recommendation action, only the trajectory with the highest advantage, identified by $i^{\star}=\arg\max_{j}A_{j}$, contributes gradients to recommendation optimization. This design concentrates recommendation learning with the most promising reasoning path, while the presence of other G1G-1 reasoning paths contributes to ample exploration in reasoning actions, thus preserving exploration and ensuring effective recommendation learning. We further provide a gradient-based explanation of how item encoding, as a core component of the model’s recommendation ability, participates in the objective and enables end-to-end optimization in Appendix F.
@@ 4.1 Setups
Following previous works [3, 2, 1, 25], we conducted experiments using real-world datasets sourced from Amazon, including CDs and Vinyl (CDs), Video Games (Games), and Musical Instruments (Instruments). Dataset statistics and preprocessing steps detailed in Appendix B. We utilized two commonly used metrics: Hit Rate (H@K) and Normalized Discounted Cumulative Gain (N@K), with cutoff $K$ set to $5$, $10$ and $20$. We adopted the full set recommendation setting, where metrics are computed over the entire item set, providing a better reflection of practical scenarios.
@@
We selected competitive baselines from various categories, including traditional sequential recommenders (GRU4Rec [36], Caser [37], SASRec [38]), LLM-based recommender (TIGER [9], BigRec [2], $D^{3}$ [3], SDPO [25], Llara [1], SPRec [39]), and reasoning augmented recommendation systems (LangPTune [16]). For fair comparison, we adapted LLaRA and SDPO by removing candidate prompts to perform constrained generation over the full item corpus, denoted as LLaRA* and SDPO* respectively. Backboned on Gemma2-2B-Instruct and Qwen2.5-3B-Instruct. More baseline and implementation details can be found in Appendix C and D, respectively.